The graph of y= -4x + 5 is shown on the next page. Check your work.
The value of x when y = -3
Not sure what the question is asking??

Answer by nyc_function(2741)

The question is asking you to replace y with -3 and solve for x.
y = -4x + 5
-3 = -4x + 5
-3 - 5 = -4x
-8 = -4x
-8/-4 = x
8/4 = x
2 = x
Done!